OpenStudy (anonymous):

1. Suppose y varies directly with x, and y=14 when x=-4. what is the value of y when x=-6? a.-21 b.10 c.18 d.21

OpenStudy (danjs):

varies directly means it is directly proportional \[\large y~ \alpha ~x\]

OpenStudy (danjs):

so, y is any constant times x y = k*x use the given point x= y= they gave to figure the k value

OpenStudy (anonymous):

thank you

OpenStudy (danjs):

then with that k you can use y=k*x, to get y when x is anything, here they want y when x=-6

OpenStudy (anonymous):

if y varies directly with x, then \[\frac{ y }{ x }=constant\]

OpenStudy (anonymous):

this means any pair of y and x gives the same value of constant... from the given pair of y and x you can find the constant and then use y = constant * x to find the missing value of y